# The Effect of a Non-universal Extinction Curve on the Wesenheit Function and Cepheid Distances

### v1

@@ Abstract
software: scipy (2020SciPy-NMeth), numpy (harris2020array), matplotlib (Hunter:2007), Jupyter notebooks (Kluyver2016jupyter), Pyphot (Fouesneau_pyphot_2025),
@@
1 Introduction
The Wesenheit [^1] Wesenheit is a German word for ”intrinsic essence” function, $W$, is an expression for the magnitude of a star that is free of interstellar extinction and reddening. It was conceptually introduced by vandenBergh1975, although the idea of constructing a reddening-free magnitude traces back to Johnson1953. The term “Wesenheit” itself was later introduced by Madore1982, who also explored its detailed properties and applications, using it to calibrate the period–luminosity relation for classical Cepheids. A comprehensive historical overview of the development and evolution of the Wesenheit function, as well as its underlying physical basis can be found in Madore2024.
While the Wesenheit function can be applied to any type of star, it has been used primarily as a luminosity indicator in the period–luminosity (PL) relation for pulsating stars, such as Cepheids, RR Lyrae, or long-period variables. This relation, expressed as $W=a\log P+b$, is commonly referred to as the period–Wesenheit (PW) relation. Its use significantly reduces the scatter relative to the standard PL relation by effectively accounting for interstellar extinction (see Fig. 6 in Soszynski2008). This is particularly important in the case of classical Cepheids, which serve as standard candles for measuring precise distances within the Milky Way and in galaxies well beyond the Local Group, making them excellent tracers of young stellar populations and key contributors to the calibration of the cosmic distance scale and the determination of the Hubble constant (Casertano2025). Therefore, the subsequent analysis will be focused on classical Cepheid variable stars.
@@ 1 Introduction
which is assumed to have a constant mean value of around $3.1$ in the Milky Way (e.g., Schultz1975; Fitzpatrick1999). However, growing evidence indicates that $R_{V}$ is not constant within the Galaxy, but it varies significantly, particularly in regions of high extinction (Fitzpatrick2007; Schlafly2016; Nataf2016; Maiz-Apellaniz2018; Fouesneau2022; Zhang2023; Zhang2025). This is especially important for classical Cepheids, which belong to the young population and are therefore confined to the Galactic disk, where the extinction is highest. The adverse consequences of a variable $R_{V}$ on Cepheid distance estimates for Galactic structure studies have already been reported by Skowron2025, and this letter aims to explore this effect in greater detail.
@@
2 The Wesenheit Function Definitions
@@
Table 1: Wesenheit index definitions used in various studies and calculated in this study using the extinction curve of Gordon2023.
@@
$W_{VI}=I-R_{VI}(V-I)$ $R_{VI}=A_{I}/E(V-I)$ $1.55$ $R_{V}=3.23$, Cardelli1989
$1.387$ $R_{V}=3.1$, Fitzpatrick1999
$1.335$ $R_{V}=3.1$, Gordon2023
$W_{JK}=K-R_{JK}(J-K)$ $R_{JK}=A_{K}/E(J-K)$ $0.679$ $R_{V}=3.1$, Cardelli1989
$0.688$ $R_{V}=3.23$, Cardelli1989
$0.553$ $R_{V}=3.1$, Gordon2023
$W_{VK}=K-R_{VK}(V-K)$ $R_{VK}=A_{K}/E(V-K)$ $0.129$ $R_{V}=3.23$, Cardelli1989
$0.090$ $R_{V}=3.1$, Gordon2023
$W_{H}=H-R_{H}(V-I)$ $R_{H}=A_{H}/E(V-I)$ $0.365$ $R_{V}=3.1$, Cardelli1989
$0.35$ $R_{V}=3.1$, Fitzpatrick1999
$0.321$ $R_{V}=3.1$, Gordon2023
$W_{G}=G-R_{G}(G_{BP}-G_{RP})$ $R_{G}=A_{G}/E(G_{BP}-G_{RP})$ $1.9$ empirical, Ripepi2019
$1.67$ $R_{V}=3.1$, Gordon2023
@@
In practice, the choice of which Wesenheit index to use is dictated by the passbands for which observational data are available. Traditionally, $B$ and $V$ bands have been employed; however, following the publication of catalogs containing over $9,500$ Cepheids in the Magellanic Clouds by the Optical Gravitational Lensing Experiment (OGLE), the $V$ and $I$ bands have become widely adopted (Soszynski2008; Soszynski2015), while additional observations from near-IR surveys have led to the use of a variety of Wesenheit indices that combine the $V$, $J$, $H$, and $K_{S}$ magnitudes. The use of classical Cepheids as standard candles, where distance measurements are based on Hubble Space Telescope (HST) observations in the optical ($V=F555W$, $I=F814W$) and near-infrared ($H=F160W$) bands, relies on the near-IR Wesenheit index, $W_{H}$, constructed from these three HST passbands (Riess2009). With the availability of all-sky data for classical Cepheids from the Gaia mission, the Wesenheit magnitude $W_{G}$ was introduced, with the three-band variant exhibiting the lowest scatter (Ripepi2019).
@@
In the following analysis, we focus on the most often used Wesenheit functions, listed in Table 1, although numerous other combinations have also been studied (e.g. Inno2016; Breuval2022; Bhardwaj2024).
3 Extinction Curve
@@
To calculate the $R$ coefficient in different bands, we need to assume a relation between interstellar dust extinction and wavelength. Commonly used $R_{V}$ dependent extinction curves in the Milky Way include those of Cardelli1989 and Fitzpatrick1999, and more recently Gordon2009; Fitzpatrick2019; Gordon2021; Decleir2022; Zhang2025, which cover various wavelength ranges. However, because these studies used different samples, methods, and wavelength ranges, combining the resulting extinction curves to cover the full ultraviolet-mid-infrared range could introduce artificial discontinuities between wavelength regions. Therefore, for the following analysis, we adopt the extinction curve of Gordon2023, which is the first $R_{V}$ dependent extinction relation valid over the full wavelength range.
@@
G23 define an extinction curve normalized to the monochromatic extinction at 550 nm. However, we stress that the wavelength dependence of their extinction curve in the optical range does not differ significantly from past formulations (e.g., Fitzpatrick1999; Fitzpatrick2019). Nevertheless, their normalization to a monochromatic extinction leads to differences that we highlight below.
@@
We used the Atlas-9 (2004; Castelli2003) atmosphere library to compute a grid of $17,603$ stellar spectra references spanning effective temperatures $T_{\rm eff}\in[2,500;8,000]$ K, surface gravities $\log g\in[-0.5;5.5]$ dex, and metallicities $M/H]\in[-5;+1]$ dex. We expand this collection by applying the G23 extinction curve on a grid of $A_{0}\in[0,10.0]$ by step of $0.5$ mag and $R_{0}\in[2.3,2.6,3.1,3.6,4.1,4.6,5.1,5.6]$. Finally, we use Pyphot (Fouesneau_pyphot_2025) to compute the photometric calculations in the relevant passbands for the present work, namely: Johnson $B$ & $V$; Cousins $I$; Gaia DR3 $G$, $G_{BP}$, $G_{RP}$ and 2MASS $J$, $H$, $K_{s}$. [^2] We took the transmission curves from the SVO-profile service: http://svo2.cab.inta-csic.es/theory/fps/; svofps2020
@@
where the coefficients have uncertainties of no more than 2%, and they remain unchanged within $T_{\rm eff}$ and $\log g$ range typical for Cepheids. While this offset remains nearly constant as $R_{0}$ is varied, the lower panel of Figure 2 shows the ratio $R_{V}$$R_{0}$ to highlight the subtle variation with respect to $R_{0}$, where the shaded region illustrates the uncertainty in the ratio from our polynomial fit. Because the parameter $R_{0}$ is calibrated on measurements of $R_{55}$, we found that the offset of $R_{0}$ with respect to $R_{V}$ (Eq. 7) is quite similar to the offset of $R_{55}$ with respect to $R_{V}$, as specified by Zhang2025, i.e. $R_{V}=1.1\,R_{55}+0.07$.
Figure 2: $R_{V}$ as a function of $R_{0}$, shown for three different types of stars. In the upper panel, the shaded region captures the spread over 0.5 $A_{0}<$ 9.5, while the points are the median values. The black solid curve is the best-fit polynomial for Cepheids (Eq. 7), while the cyan line is the proposed relation between $R_{V}$ and $R_{55}$ of Zhang2025. The grey dotted line indicates the identity relation. The lower panel shows the ratio $R_{V}$$R_{0}$ with the spread capturing the uncertainty in the polynomial fit.
4 Discussion
@@
Figure 3: The dependence of the $R$ coefficient, used in the definition of a given Wesenheit index, on $R_{V}$, as a function of the color excess $E$ for various passbands, calculated using a typical Cepheid source spectra. Colors mark different $R_{V}$ values, according to the legend. Solid lines assume the G23 extinction curve, while the light-green dashed line is the Fitzpatrick1999 extinction curve, but only for $R_{V}=3.1$. Black dashed lines show constant $R$ coefficients adopted in various studies, and the grey dotted line represents reddening values typical for Milky Way Cepheids.
Recently Zhang2025 measured the $R_{V}$ parameter within the Milky Way and the Magellanic Clouds using low-resolution optical spectra from Gaia for $\sim 130$ million stars. The resulting three-dimensional maps provide $R_{V}$ values as a function of distance, enabling more accurate extinction corrections than those obtained by assuming a constant $R_{V}$ value. The mean differential $R_{V}$ value in the Galaxy is $3.11$ with a $1\sigma$ range from $2.60$ to $3.88$, and the average integrated $R_{V}$ is $3.08$ with a $1\sigma$ range from $2.86$ to $3.28$, although values as low as $2.00$ and as high as $8.00$ are observed. Both average estimates are consistent with $R_{V}=3.1$, as found in many previous studies of the Milky Way. This section shows how varying extinction curve, as parameterized by $R_{V}$, affects the Wesenheit magnitudes and distances based on PW relations. That is, we fix the extinction curve using a specific value of the parameter $R_{0}$, but in this section translate these $R_{0}$ values to the more commonly used $R_{V}$ parameter used in the past to parameterize extinction curves, using Eq. 7 specific to Cepheids.
4.1 The impact of $R_{V}$ on the $R$ coefficient
The Wesenheit index parameter $R$ quantifies the ratio of total to selective extinction for a given combination of passbands (Eqns. 3,5) and therefore varies across specific definitions of the Wesenheit index. Figure 3 shows the $R$ coefficient as a function of the color excess $E$, where each panel corresponds to a different definition of $W$ and $R$, listed in Table 1. The color coding in each panel represents distinct $R_{V}$ values, and the grey dotted line marks the reddening values typical for Milky Way Cepheids. Light-green dashed lines represent $R$, when instead of G23, the Fitzpatrick1999 extinction curve for $R_{V}=3.1$ is used, and black dashed lines show constant values used in the literature.
@@ 4.1 The impact of $R_{V}$ on the $R$ coefficient
However, a more important conclusion from Figure 3 is that the value of $R$ varies significantly with $R_{V}$. For example, in the first panel of this figure, we see that for a typical Cepheid reddening (indicated by the gray dotted line), the value of $R_{VI}$ changes from $1.05$ to $1.6$ for $R_{V}$ varying from 2.6 to 3.6, which is a typical $R_{V}$ range in the Milky Way. In Section 4.2 we will show how this impacts the Wesenheit magnitude and the resulting distance determinations.
@@
4.2 The impact of $R_{V}$ on the Wesenheit magnitude and distance
Following the results of Zhang2025, we adopt the mean value of $R_{V}=3.1\pm 0.22$. Then $R_{V}=2.6$ and $R_{V}=3.6$, which represent a typical $R_{V}$ range in the Milky Way, are a good approximation of the $\pm 2\sigma$ range around the mean of the integrated $R_{V}$ [^4] Zhang2025 provide maps for both differential, and integrated along a line-of-sight to a specific distance, $R_{55}$, from which the integrated $R_{V}$ (Eq. 1) can be derived. (see their Figure S8). In that case, the relative difference in the $R$ parameter within $R_{V}=3.1\pm 0.22$, estimated from Figure 4, is roughly 8.5%, 10%, 12%, 21%, and 21.5% for $R_{G}$, $R_{VI}$, $R_{JK}$, $R_{H}$, and $R_{VK}$, respectively.

### v2

@@ Abstract
software: scipy (Virtanen et al., 2020), numpy (Harris et al., 2020), matplotlib (Hunter, 2007), Jupyter notebooks (Kluyver et al., 2016), Pyphot (Fouesneau, 2025),
@@
I Introduction
The Wesenheit [^1] Wesenheit is a German word for ”intrinsic essence” function, $W$, is an expression for the magnitude of a star that is free of interstellar extinction and reddening. It was conceptually introduced by van den Bergh (1975), although the idea of constructing a reddening-free magnitude traces back to Johnson and Morgan (1953). The term “Wesenheit” itself was later introduced by Madore (1982), who also explored its detailed properties and applications, using it to calibrate the period–luminosity relation for classical Cepheids. A comprehensive historical overview of the development and evolution of the Wesenheit function, as well as its underlying physical basis can be found in Madore and Freedman (2024).
While the Wesenheit function can be applied to any type of star, it has been used primarily as a luminosity indicator in the period–luminosity (PL) relation for pulsating stars, such as Cepheids, RR Lyrae, or long-period variables. This relation, expressed as $W=a\log P+b$, is commonly referred to as the period–Wesenheit (PW) relation. Its use significantly reduces the scatter relative to the standard PL relation by effectively accounting for interstellar extinction (see Fig. 6 in Soszyński et al., 2008). This is particularly important in the case of classical Cepheids, which serve as standard candles for measuring precise distances within the Milky Way and in galaxies well beyond the Local Group, making them excellent tracers of young stellar populations and key contributors to the calibration of the cosmic distance scale and the determination of the Hubble constant (H0DN Collaboration et al., 2025). Therefore, the subsequent analysis will be focused on classical Cepheid variable stars.
@@ I Introduction
which is assumed to have a constant mean value of around $3.1$ in the Milky Way (e.g., Schultz and Wiemer, 1975; Fitzpatrick, 1999). However, growing evidence indicates that $R_{V}$ is not constant within the Galaxy, but it varies significantly, particularly in regions of high extinction (Fitzpatrick and Massa, 2007; Schlafly et al., 2016; Nataf et al., 2016; Maíz Apellániz and Barbá, 2018; Fouesneau et al., 2022; Zhang et al., 2023; Zhang and Green, 2025). This is especially important for classical Cepheids, which belong to the young population and are therefore confined to the Galactic disk, where the extinction is highest. The adverse consequences of a variable $R_{V}$ on Cepheid distance estimates for Galactic structure studies have already been reported by Skowron et al. (2025), and this letter aims to explore this effect in greater detail.
@@
II The Wesenheit Function Definitions
@@
Table 1: Wesenheit index definitions used in various studies and calculated in this study using the extinction curve of Gordon et al. (2023).
@@
$W_{VI}=I-R_{VI}(V-I)$ $R_{VI}=A_{I}/E(V-I)$ $1.55$ $R_{V}=3.23$, Cardelli et al. (1989)
$1.387$ $R_{V}=3.1$, Fitzpatrick (1999)
$1.335$ $R_{V}=3.1$, Gordon et al. (2023)
$W_{JK}=K-R_{JK}(J-K)$ $R_{JK}=A_{K}/E(J-K)$ $0.679$ $R_{V}=3.1$, Cardelli et al. (1989)
$0.688$ $R_{V}=3.23$, Cardelli et al. (1989)
$0.553$ $R_{V}=3.1$, Gordon et al. (2023)
$W_{VK}=K-R_{VK}(V-K)$ $R_{VK}=A_{K}/E(V-K)$ $0.129$ $R_{V}=3.23$, Cardelli et al. (1989)
$0.090$ $R_{V}=3.1$, Gordon et al. (2023)
$W_{H}=H-R_{H}(V-I)$ $R_{H}=A_{H}/E(V-I)$ $0.365$ $R_{V}=3.1$, Cardelli et al. (1989)
$0.35$ $R_{V}=3.1$, Fitzpatrick (1999)
$0.321$ $R_{V}=3.1$, Gordon et al. (2023)
$W_{G}=G-R_{G}(G_{BP}-G_{RP})$ $R_{G}=A_{G}/E(G_{BP}-G_{RP})$ $1.9$ empirical, Ripepi et al. (2019)
$1.67$ $R_{V}=3.1$, Gordon et al. (2023)
@@
In practice, the choice of which Wesenheit index to use is dictated by the passbands for which observational data are available. Traditionally, $B$ and $V$ bands have been employed; however, following the publication of catalogs containing over $9,500$ Cepheids in the Magellanic Clouds by the Optical Gravitational Lensing Experiment (OGLE), the $V$ and $I$ bands have become widely adopted (Soszyński et al., 2008, 2015), while additional observations from near-IR surveys have led to the use of a variety of Wesenheit indices that combine the $V$, $J$, $H$, and $K_{S}$ magnitudes. The use of classical Cepheids as standard candles, where distance measurements are based on Hubble Space Telescope (HST) observations in the optical ($V=F555W$, $I=F814W$) and near-infrared ($H=F160W$) bands, relies on the near-IR Wesenheit index, $W_{H}$, constructed from these three HST passbands (Riess et al., 2009). With the availability of all-sky data for classical Cepheids from the Gaia mission, the Wesenheit magnitude $W_{G}$ was introduced, with the three-band variant exhibiting the lowest scatter (Ripepi et al., 2019).
@@
In the following analysis, we focus on the most often used Wesenheit functions, listed in Table 1, although numerous other combinations have also been studied (e.g. Inno et al., 2016; Breuval et al., 2022; Bhardwaj et al., 2024).
III Extinction Curve
@@
To calculate the $R$ coefficient in different bands, we need to assume a relation between interstellar dust extinction and wavelength. Commonly used $R_{V}$ dependent extinction curves in the Milky Way include those of Cardelli et al. (1989) and Fitzpatrick (1999), and more recently Gordon et al. (2009); Fitzpatrick et al. (2019); Gordon et al. (2021); Decleir et al. (2022); Zhang and Green (2025), which cover various wavelength ranges. However, because these studies used different samples, methods, and wavelength ranges, combining the resulting extinction curves to cover the full ultraviolet-mid-infrared range could introduce artificial discontinuities between wavelength regions. Therefore, for the following analysis, we adopt the extinction curve of Gordon et al. (2023, hereafter G23), which is the first $R_{V}$ dependent extinction relation valid over the full wavelength range.
@@
G23 define an extinction curve normalized to the monochromatic extinction at 550 nm. However, we stress that the wavelength dependence of their extinction curve in the optical range does not differ significantly from past formulations (e.g., Fitzpatrick 1999; Fitzpatrick et al. 2019). Nevertheless, their normalization to a monochromatic extinction leads to differences that we highlight below.
@@
We used the Atlas-9 (2004; Castelli and Kurucz 2003) atmosphere library to compute a grid of $17,603$ stellar spectra references spanning effective temperatures $T_{\rm eff}\in[2,500;8,000]$ K, surface gravities $\log g\in[-0.5;5.5]$ dex, and metallicities $M/H]\in[-5;+1]$ dex. We expand this collection by applying the G23 extinction curve on a grid of $A_{0}\in[0,10.0]$ by step of $0.5$ mag and $R_{0}\in[2.3,2.6,3.1,3.6,4.1,4.6,5.1,5.6]$. Finally, we use Pyphot (Fouesneau, 2025) to compute the photometric calculations in the relevant passbands for the present work, namely: Johnson $B$ & $V$; Cousins $I$; Gaia DR3 $G$, $G_{BP}$, $G_{RP}$ and 2MASS $J$, $H$, $K_{S}$. [^2] We took the transmission curves from the SVO-profile service: http://svo2.cab.inta-csic.es/theory/fps/; Rodrigo and Solano (2020)
@@
where the coefficients have uncertainties of no more than 2%, and they remain unchanged within $T_{\rm eff}$ and $\log g$ range typical for Cepheids. While this offset remains nearly constant as $R_{0}$ is varied, the lower panel of Figure 2 shows the ratio $R_{V}$$R_{0}$ to highlight the subtle variation with respect to $R_{0}$, where the shaded region illustrates the uncertainty in the ratio from our polynomial fit. Because the parameter $R_{0}$ is calibrated on measurements of $R_{55}$, we found that the offset of $R_{0}$ with respect to $R_{V}$ (Eq. 7) is quite similar to the offset of $R_{55}$ with respect to $R_{V}$, as specified by Zhang and Green (2025), i.e. $R_{V}=1.1\,R_{55}+0.07$.
Figure 2: $R_{V}$ as a function of $R_{0}$, shown for three different types of stars. In the upper panel, the shaded region captures the spread over 0.5 $A_{0}<$ 9.5, while the points are the median values. The black solid curve is the best-fit polynomial for Cepheids (Eq. 7), while the cyan line is the proposed relation between $R_{V}$ and $R_{55}$ of Zhang and Green (2025). The grey dotted line indicates the identity relation. The lower panel shows the ratio $R_{V}$$R_{0}$ with the spread capturing the uncertainty in the polynomial fit.
IV Discussion
@@
Figure 3: The dependence of the $R$ coefficient, used in the definition of a given Wesenheit index, on $R_{V}$, as a function of the color excess $E$ for various passbands, calculated using a typical Cepheid source spectra. Colors mark different $R_{V}$ values, according to the legend. Solid lines assume the G23 extinction curve, while the light-green dashed line is the Fitzpatrick (1999) extinction curve, but only for $R_{V}=3.1$. Black dashed lines show constant $R$ coefficients adopted in various studies, and the grey dotted line represents reddening values typical for Milky Way Cepheids.
Recently Zhang and Green (2025) measured the $R_{V}$ parameter within the Milky Way and the Magellanic Clouds using low-resolution optical spectra from Gaia for $\sim 130$ million stars. The resulting three-dimensional maps provide $R_{V}$ values as a function of distance, enabling more accurate extinction corrections than those obtained by assuming a constant $R_{V}$ value. The mean differential $R_{V}$ value in the Galaxy is $3.11$ with a $1\sigma$ range from $2.60$ to $3.88$, and the average integrated $R_{V}$ is $3.08$ with a $1\sigma$ range from $2.86$ to $3.28$, although values as low as $2.00$ and as high as $8.00$ are observed. Both average estimates are consistent with $R_{V}=3.1$, as found in many previous studies of the Milky Way. This section shows how varying extinction curve, as parameterized by $R_{V}$, affects the Wesenheit magnitudes and distances based on PW relations. That is, we fix the extinction curve using a specific value of the parameter $R_{0}$, but in this section translate these $R_{0}$ values to the more commonly used $R_{V}$ parameter used in the past to parameterize extinction curves, using Eq. 7 specific to Cepheids.
IV.1 The impact of $R_{V}$ on the $R$ coefficient
The Wesenheit index parameter $R$ quantifies the ratio of total to selective extinction for a given combination of passbands (Eqns. 3,5) and therefore varies across specific definitions of the Wesenheit index. Figure 3 shows the $R$ coefficient as a function of the color excess $E$, where each panel corresponds to a different definition of $W$ and $R$, listed in Table 1. The color coding in each panel represents distinct $R_{V}$ values, and the grey dotted line marks the reddening values typical for Milky Way Cepheids. Light-green dashed lines represent $R$, when instead of G23, the Fitzpatrick (1999) extinction curve for $R_{V}=3.1$ is used, and black dashed lines show constant values used in the literature.
@@ IV.1 The impact of $R_{V}$ on the $R$ coefficient
However, a more important conclusion from Figure 3 is that the value of $R$ varies significantly with $R_{V}$. For example, in the first panel of this figure, we see that for a typical Cepheid reddening (indicated by the gray dotted line), the value of $R_{VI}$ changes from $1.05$ to $1.6$ for $R_{V}$ varying from 2.6 to 3.6, which is a typical $R_{V}$ range in the Milky Way. In Section IV.2 we will show how this impacts the Wesenheit magnitude and the resulting distance determinations.
@@
IV.2 The impact of $R_{V}$ on the Wesenheit magnitude and distance
Following the results of Zhang and Green (2025), we adopt the mean value of $R_{V}=3.1\pm 0.22$. Then $R_{V}=2.6$ and $R_{V}=3.6$, which represent a typical $R_{V}$ range in the Milky Way, are a good approximation of the $\pm 2\sigma$ range around the mean of the integrated $R_{V}$ [^4] Zhang and Green (2025) provide maps for both differential, and integrated along a line-of-sight to a specific distance, $R_{55}$, from which the integrated $R_{V}$ (Eq. 1) can be derived. (see their Figure S8). In that case, the relative difference in the $R$ parameter within $R_{V}=3.1\pm 0.22$, estimated from Figure 4, is roughly 8.5%, 10%, 12%, 21%, and 21.5% for $R_{G}$, $R_{VI}$, $R_{JK}$, $R_{H}$, and $R_{VK}$, respectively.